OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226. -334 -274 -284 -346

OpenStudy (michele_laino):

hint: we can write these equations: \[\Large \begin{gathered} {a_{22}} = {a_1} + 21d \hfill \\ {a_{31}} = {a_1} + 30d \hfill \\ \end{gathered} \] since the general formula is: \[\Large {a_n} = {a_1} + \left( {n - 1} \right)d\] where d is the constant of your sequence

OpenStudy (michele_laino):

using your data we can rewrite the first equation as follows: \[\Large - 226 = 26 + 21d\] please solve that equation for d

OpenStudy (anonymous):

-252=21d d=-12

OpenStudy (anonymous):

is that right

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, substituting that value of d into the second equation, we get: \[\Large {a_{31}} = 26 + 30 \times \left( { - 12} \right) = ...?\]

OpenStudy (anonymous):

-344

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

Given the functions f(n ) = 11 and g(n ) = -2(n - 1), combine them to create an arithmetic sequence, an, and solve for the 31st term. an = 11 - 2(n - 1); a31 = -49 an = 11 - 2(n - 1); a31 = -51 an = 11 + 2(n - 1); a31 = 71 an = 11 + 2(n - 1); a31 = 73

OpenStudy (anonymous):

can u help me with this one

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

for example, let's consider the first option: we have: \[\Large {a_n} = 11 - 2\left( {n - 1} \right)\] so for n=31, we can rewrite that equationas follows: \[\Large {a_{31}} = 11 - 2 \times \left( {31 - 1} \right) = ...?\] please continue

OpenStudy (anonymous):

a31= -49 right?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Given an arithmetic sequence in the table below, create the explicit formula and list any restrictions to the domain. n an 1 40 2 47 3 54

OpenStudy (anonymous):

can u help me with this one

OpenStudy (anonymous):

Those are the options: an = 40 + 7(n - 1) where n ≥ 40 an = 40 + 7(n - 1) where n ≥ 1 an = 40 - 7(n - 1) where n ≥ 40 an = 40 - 7(n - 1) where n ≥ 1

OpenStudy (michele_laino):

Please wait: also the third option is correct, since we can write this: \[\Large {a_{31}} = 11 + 2 \times \left( {31 - 1} \right) = ...?\]

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

a31=11+2×(31−1)= 71

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

71 is not an answer option tho

OpenStudy (anonymous):

Oh yes it is nevermind

OpenStudy (anonymous):

Which one is the right one? How can I know?

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (anonymous):

Do I just make a guess?

OpenStudy (michele_laino):

maybe the first one, since the first option contains both f(n) and g(n), whereas the third option contains f(n) and -g(n)

OpenStudy (anonymous):

Okay thanks!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

can u help me with thise one now: Given an arithmetic sequence in the table below, create the explicit formula and list any restrictions to the domain. n an 1 40 2 47 3 54 Those are the options: an = 40 + 7(n - 1) where n ≥ 40 an = 40 + 7(n - 1) where n ≥ 1 an = 40 - 7(n - 1) where n ≥ 40 an = 40 - 7(n - 1) where n ≥ 1

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

the constant of your sequence is: \[d = 47 - 40 = 54 - 47 = ...?\]

OpenStudy (anonymous):

d=7

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

:)

OpenStudy (michele_laino):

so, since the general formula, is: \[\Large {a_n} = {a_1} + \left( {n - 1} \right)d\] replace a_1 with 40 and d with 7, what do you get?

OpenStudy (anonymous):

an=40+(n−1)7

OpenStudy (anonymous):

whats in the n spot?

OpenStudy (michele_laino):

n is the number of terms, it is a natural number, more precisely n-1 is the number of terms of the sequence which precede a_n

OpenStudy (anonymous):

Okay so is the answer the second one

OpenStudy (michele_laino):

yes! since we start to count from n=1

OpenStudy (anonymous):

THANKS!

OpenStudy (michele_laino):

:)